(Bond valuation​)

Bellingham bonds have an annual coupon rate of 13 percent and a par value of $1,000 and will mature in 5 years. If you require a return of 12 ​percent, what price would you be willing to pay for the​ bond? What happens if you pay more for the​ bond? What happens if you pay less for the​ bond?

a. The price you would be willing to pay for the bond is

​ (Round to the nearest​ cent.)
Answer #1

Answer:

Par value = $1,000

Annual coupon = 1000 * 13% = $130

Time to maturity = 5 years

Required rate of return = 12%

Hence:

The price you would be willing to pay for the bond is = PV (rate, nper, pmt, fv, type) = PV(12%, 5, -130, -1000, 0) =$1036.05

The price you would be willing to pay for the bond is = $1,036.05

If you pay more for the bond you will be earning lesser rate of return than the required rate.

If you pay less for the bond you will be earning higher rate of return than the required rate.